OpenStudy (anonymous):

the frequency of a certain wave is 500Hz and its speed is 340 ms^-1 what is the phase difference between the motions of the two points on the wave 0.17m apart?

OpenStudy (kropot72):

\[wavelength=\frac{speed}{frequency}=\frac{340}{500}\ m\] One wavelength occupies one cycle. One cycle takes 360 degrees. Therefore the required phase difference is given by \[360\times \frac{0.17}{wavelength}\]

OpenStudy (anonymous):

the answer is pie upon 2 rad . how come?

OpenStudy (kropot72):

\[\frac{\pi}{2}\ radians=90\ degrees\]

OpenStudy (anonymous):

Do you need help solving the entire thing?

OpenStudy (kropot72):

If the answer is required in radians the solution is given by \[2 \pi \times \frac{0.17}{0.68}\]

OpenStudy (anonymous):

@blackilluminati yes

OpenStudy (anonymous):

frequency= 500hz Speed= 340ms-1 Relevant equations x= d x wavelength/ a Wavelength= Speed/frequency 340/500= 0.68m Two points will be out of phase by 360* (*=Degree) 0.68m= 360* therefore what is the angle for 0.17m=? You cross multiply 0.17 x 360/ 0.68= 90* There you go (: